Find the future value of the following investment.

Principal

Nominal Rate

Frequency of Conversion

Time

​$5530.00

6.9​%

monthly

5 years
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=5530*(1+0.069/12)^(12*5)

=5530*1.41059544

=$7800.59(Approx).